# SqueezeAttention: 2D Management of KV-Cache in LLM Inference via Layer-wise Optimal Budget

### v1

@@ 5.2 End-to-End Result
5.3 Memory Consumption
Table 2 demonstrates that in all three settings, by achieving the same model accuracy, SqueezeAttention consumes the least amount of KV-cache budgets compared with the algorithms that only compress from the sequence’s dimension. In some cases, it only takes one-third of the cache budget of H2O algorithm. Subsequently, we employ PYTORCH PROFILER to evaluate the diminished memory usage of generating one token during the inference (w/o memory usage of model weights). Figure 4 shows that SqueezeAttention can save 70% to 80% of memory usage per token compared with Full Cache method, and 25% to 66% of memory usage compared with baseline algorithms.
5.4 Throughput of Token Generation
Since SqueezeAttention manages to save the memory cost of inference, as shown in the previous sections. Now we want to explore how these memory reductions can be interpreted into improvements of token throughput. We choose 2 models, i.e., Mistral-7B and Llama-70B, to represent models in different sizes. With a fixed content length, we increase the batch size from 1 to 224 for Mistral-7B, and 1 to 64 for Llama2. For each task, we select the best setting for the SqueezeAttention, that is, Sliding Window for Mistral-7B and StreamingLLM for Llama2-70B.
Table 3 shows the token throughput on 8 A100-40GB GPUs. With the same batch size, SqueezeAttention can enhance throughput by up to 2.2$\times$ for Mistral-7B and 1.4$\times$ for Llama2-70B compared to the Full Cache. Besides, SqueezeAttention also enables batch sizes up to 224 and 64 for two models, which would cause out-of-memory for the Full Cache method.
@@ 6 Conclusion
In this paper, we propose a 2D KV-cache compression algorithm called SqueezeAttention. By tracking the cosine similarity of each attention layer, we found that layers in different positions have distinct degrees of importance regarding the output embedding. Inspired by this observation, SqueezeAttention reallocates the KV-cache budgets over attention layers to further reduce the memory cost of inference. Experiments over a wide range of models and tasks show that SqueezeAttention can achieve better model accuracies with lower memory consumption compared with SOAT algorithms that only compress KV-cache from one dimension.

### v2

@@ 5.4 Throughput of Token Generation
5.5 Overhead of the Algorithm
The computational overhead of SqueezeAttention comes from two operations: Cosine similarity and Kmeans. The execution of all these computations only happens during the prefilling phase. Therefore, the cost of SqueezeAttention is a one-time price, which is much more cost-efficient than those algorithms (like H2O) that require extra calculations in each iteration of token generation.
We profiled the entire prefilling phase of Mistral-7B with/without SqueezeAttention to compare the wall-clock time. The prompt length is up to 8k tokens. As shown in Table 4, the SqueezeAttention only caused 6.3% increasement of prefilling time. Note that if we take the decoding time into consideration, this ratio is going to be much smaller, depending on the actual number of tokens generated. Therefore, the overhead of SqueezeAttention is basically neglectable. The more detailed overhead of SqueezeAttention can be found in 5.
Table 4: Overhead of SqueezeAttention (Prefilling Time in seconds on one Nvidia A100-40GB)
Model w/o SqueezeAttention w/ SqueezeAttention Overhead Ratio
Mistral-7B 0.636 0.676 6.3%
5.6 Limitations and Broader Impacts
SqueezeAttention works jointly with a sequence-wise sparsification policy, so the assumption we make is that for a given model and dataset, there exists a sequence-wise KV-cache eviction policy that won’t hurt the model accuracy under a certain tolerance. However, the generalizability of these sequence-wise algorithms is still an active research topic. If this assumption cannot be met, SqueezeAttention might not work as expected.
The positive social impacts include reducing the energy cost and carbon emissions of LLM inference. Besides, it can optimize the user experience of LLM applications. Since the proposed algorithm could accelerate the LLM inference, the potential negative impact is that it might exaggerate the improper usage of LLM, like generating harmful information.
@@ 6 Conclusion
In this paper, we propose a 2D KV-cache compression algorithm called SqueezeAttention. By tracking the cosine similarity of each attention layer, we found that layers in different positions have distinct degrees of importance regarding the output embedding. Inspired by this observation, SqueezeAttention reallocates the KV-cache budgets over attention layers to further reduce the memory cost of inference. Experiments over a wide range of models and tasks show that SqueezeAttention can achieve better model accuracies with lower memory consumption compared with state-of-the-art algorithms that only compress KV-cache in a sequence-wise way.
Appendix A Appendix
A.1 detailed overhead of squeezeattention
In this section, we further evaluate the overhead introduced by SqueezeAttention, we broke down the time taken for two primary operations: cosine similarity computation and K-means clustering. The experiment setup follows the same setup as 5.5. This experiment was conducted using a single Nvidia A100-40GB GPU with prompt lengths of up to 8k tokens.
Table 5: detailed onverhead
Cosine similarity K-means Total time
0.00068s 0.001s 0.02276s
The operation of computing cosine similarity involves calculating between two arrays of size 8000×4096, repeated 32 times (since Mistral has 32 layers). Additionally, K-means clusters 32 numbers into 3 classes. Therefore, the total time consumption can be calculated as 0.00068×32+0.001=0.02276 seconds. It is noteworthy that this overhead is incurred only once, regardless of the number of tokens processed.
A.2 the function of $p$
The hyperparameter $p$ is crucial in the SqueezeAttention mechanism, as it directly determines the final allocation of the KV cache. To illustrate this with a concrete example: suppose we have a model with 32 layers, where 18 layers are deemed important and the remaining 14 are considered less important. Each layer initially has a budget of 1000 tokens. If we set $p$ to 0.3, we will take 70% of the budget from the less important layers and redistribute it equally among the important layers.
Specifically, the budget for the less important layers will be reduced to 1000×0.3=300. Meanwhile, the budget for the important layers will be calculated as follows: (1000×18+1000×0.7×14)/18=1544. In this way, the total budget remains unchanged.
To evaluate the sensitivity of the hyperparameter $p$, we tested the Mistral-7B model on the Samsum dataset with $p$ values ranging from 0.1 to 1.0. The total KV budget was set to 20% of the prompt length. The results are presented in the table below:
Table 6: The precision changing with $p$
ROUGE-L 15.7 34.01 37.26 37.69 27.41 26.29 11.11 10.48 8.72 9.07
p 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
100% means we do not alter the structure of the model’s KV cache. As $p$ decreases, more KV budget will be transferred to other layers, which can yield better results up to a certain point. When $p$ decreases to 10% or even lower, the performance decreases as the less important layers’ KV budget becomes severely inadequate. We can clearly observe an accuracy improvement when adjusting $p$ while keeping the overall budget unchanged. This demonstrates the impact of $p$ on model performance, highlighting its importance in optimizing KV budget allocation.
A.3 layer importance across different task
We conducted an extra experiment using two models and various datasets to determine whether the importance of different layers is an intrinsic property of the model.
Table 7: Mistral-7B
Dataset Samsum TriviaQA LCC
Important 17 18 19
Unimportant 15 14 13
Table 8: LLama2-70B
Dataset Xsum Samsum LCC
Important 17 21 18
Unimportant 63 59 62
Tabel8 displays the distribution of important layers for the Mistral model across three different datasets: Samsum (Few shot), TriviaQA (Single-document QA), and LCC (Code, Python/C#/Java) and table 8 shows the distribution of important layers for the LLama2-70B model across three different datasets: Xsum (Summarization), Samsum (Few shot), and LCC (Code, Python/C#/Java).
From these tables, we can observe that there is a rough pattern regarding the layer’s group with task-specific fluctuations. We believe there exist some task-sensitive layers that may be classified into different groups with different tasks. Similarly, there are also some layers that are always important / unimportant. A detailed analysis of this phenomenon could be an interesting extension of this work. However, we would still recommend an adaptive way since it can precisely capture the importance of layers.
A.4 throughput comparison between best baseline algorithms and SqueezeAttention
Table 9: Generation throughput (token / s) on eight A100 GPUs of Mistral-7B and LLama2-7B with SqueezeAttention and best baseline algorithms. To maintain the model accuracy, SqueezeAttention uses 20% of the cache budget for Mistral-7B and 40% of the cache budget for LLama2-7B while Sliding Window uses 30% of the cache budget for Mistral-7B and StreamingLLM use 60% of the cache budget for LLama2-7B. ”OOM” means out-of-memory.
Model Size prompt len + gen len Algorithm Batch size
1 32 64 128 224
Mistral 7B 512 + 1024 SqueezeAttention 20.5 496.5 682.7 824.4 892.5
Sliding Window 20.6 404.5 512.2 587.8 OOM
Model Size prompt len + gen len Algorithm Batch size
1 32 64 128 256
LLama2 7B 512 + 1024 SqueezeAttention 20.0 143.0 150.4 144.9 OOM
StreamLLM 20.4 113.7 102.4 OOM OOM
We also conducted experiments to compare throughput of SqueezeAttention with best baseline under a set of batch sizes. Both experiments used an input length of 512 and an output length of 1024. We choose the compression hyperparameters for each algorithm such that they could all achieve the best mode accuracy. The result show that our algorithm can obviously increase the throughput compared with those SOTA algorithms that only compress KV-cache from the sequence’s dimension.